# supplementary_materials

### v1

@@ D.2 Spiral X-DW
[…]
@@
$=-\sum_{\mathbf{q},\mathbf{k}}\vec{f}_{\mathbf{q},\mathbf{k}}^{\ {\dagger}}\Big{[}\vec{\delta\chi}({-\mathbf{q}})\cdot\vec{\sigma}\Big{]}\vec{f}_{\mathbf{q},\mathbf{k}},$ (31)
where we have used the fact that $\delta\chi_{r}=\delta\chi_{x}(\mathbf{r})+i\delta\chi_{y}(\mathbf{r})$; this implies that $[\delta\chi_{x/y}(\mathbf{q})]^{*}=\delta\chi_{x/y}(-\mathbf{q})$, and $S_{x}(\mathbf{q})\equiv\sum_{\mathbf{k}}\left(f^{\dagger}_{v,\mathbf{k}+\mathbf{q}+\mathbf{Q}}f_{c,\mathbf{k}}+f^{\dagger}_{c,\mathbf{k}}f_{v,\mathbf{k}+\mathbf{q}+\mathbf{Q}}\right)$, $S_{y}(\mathbf{q})\equiv-i\sum_{\mathbf{k}}\left(-f^{\dagger}_{v,\mathbf{k}+\mathbf{q}+\mathbf{Q}}f_{c,\mathbf{k}}+f^{\dagger}_{c,\mathbf{k}}f_{v,\mathbf{k}+\mathbf{q}+\mathbf{Q}}\right)$, and $\vec{f}_{\mathbf{q},\mathbf{k}}=(f_{c,\mathbf{k}}\ f_{v,\mathbf{k}+\mathbf{q}+\mathbf{Q}})^{\text{T}}$ where we express the fermionic operators in the basis of the mean-field Hamiltonian Eq. 12.
@@
$H_{fluc.}=\frac{A}{2}\sum_{\mathbf{q}}\delta\chi_{i}(-\mathbf{q})\left(\frac{2}{U_{0}}\delta_{ij}-\Pi_{ij}(\mathbf{q},\Omega=0)\right)\delta\chi_{j}(\mathbf{q}).$ (32)
@@
$\Pi_{ij}(\mathbf{q},\Omega)=-\int\frac{d^{2}kd\omega}{(2\pi)^{3}}\text{Tr}\left(\sigma_{i}G(\mathbf{k}+\mathbf{q},\omega+\Omega)\sigma_{j}G(\mathbf{k},\omega)\right)$ (33)
@@
$H_{fluc.}=\sum_{\mathbf{q}}\begin{pmatrix}\delta\chi_{x}(-\mathbf{q})&\delta\chi_{y}(-\mathbf{q})\end{pmatrix}\begin{pmatrix}\rho^{x}_{xx}q_{x}^{2}+\rho^{y}_{xx}q_{y}^{2}+m_{xx}&is_{xy}q_{x}\\ -is_{xy}q_{x}&\rho^{x}_{yy}q_{x}^{2}+\rho^{y}_{yy}q_{y}^{2}\end{pmatrix}\begin{pmatrix}\delta\chi_{x}(\mathbf{q})\\ \delta\chi_{y}(\mathbf{q})\end{pmatrix}$ (36)
$\rho_{xx/yy}$, $s_{xy}$ and $m_{xx}$ are fit parameters obtained numerically. The form of Eq. 36 follows from the even (odd)-in-momentum property of the diagonal (off-diagonal) elements of $\Pi_{\alpha\beta}(\mathbf{q})$.
@@
$\sin X(\mathbf{q})$ $=\frac{-s_{xy}q_{x}}{\sqrt{\left((\rho^{x}_{xx}-\rho^{x}_{yy})q_{x}^{2}+(\rho^{y}_{xx}-\rho^{y}_{yy})q_{y}^{2}+m_{xx}\right)^{2}/4+s_{xy}^{2}q_{x}^{2}}}$ (37)
Our next step is to calculate the interactions between the collective modes and the low-energy fermionic quasiparticles, as done for the collear case. We shall focus on the Goldstone fluctuations in the following, since the gapped collective mode would mediate relatively weaker interactions between the fermions away from the quantum critical line. Since we are considering hole doping, the Fermi surface is formed by the lower band $f_{-}$, which is obtained by diagonalizing the mean-field Hamiltonian 12 within the Hubbard approximation to yield eigenmodes,
$f_{-,\mathbf{k}}=-\sin\left(\frac{\theta(\mathbf{k})}{2}\right)f_{c,\mathbf{k}}+\cos\left(\frac{\theta(\mathbf{k})}{2}\right)f_{v,\mathbf{k}+\mathbf{Q}},$ (38)
where $\sin\theta(\mathbf{k})=-\frac{\chi}{E_{\mathbf{k}}}$. We can thus rewrite Eq. 31 in terms of the Goldstone mode and lower-band fermion operators to obtain the spin-independent) fermion-Goldstone mode interaction $\alpha$, where the form factor is:
$\alpha(\mathbf{k},\mathbf{q})=\cos\left(\frac{X(\mathbf{q})}{2}\right)\sin\left(\frac{\theta(\mathbf{k}+\mathbf{q})-\theta(\mathbf{k})}{2}\right)+\sin\left(\frac{X(\mathbf{q})}{2}\right)\sin\left(\frac{\theta(\mathbf{k}+\mathbf{q})+\theta(\mathbf{k})}{2}\right).$ (39)
Finally, integrating out the Goldstone mode, we obtain an effective four-fermion interaction term identical to the one derived for the collinear case in Eq.25. Fig.10 (right panel) plots the interaction $V_{G}$ as $l_{x}\rightarrow k_{x}$, $l_{y}=k_{y}$. The interaction is attractive in a region of momentum space. However, since the spiral X-DW state breaks $C_{2z}$, states at momenta $\mathbf{k}$ and $-\mathbf{k}$ are not degenerate, and it would require strong attractive interactions for them to overcome the energy difference and pair. While such a superconducting instability can occur in principle, our mean-field calculations for superconductivity do not yield superconducting solutions in the parameter space explored in this work.
Figure 10: Spiral X-DW. Left panel: Fermi surface of the spiral X-DW with yellow marking occupied states. Right panel: Goldstone mode mediated interaction matrix elements. Dashed lines plot the locus of points satisfying $\xi_{c,\mathbf{k}}=\xi_{v,\mathbf{k}\pm\mathbf{Q}}$ (curve to the left/right). The parameters used are $n=-0.7\times 10^{12}$ cm-2, $V_{b}=14.4$ meV and a $59\times 59$ momentum mesh.
Appendix E Superconductivity Mediated by Critical Fluctuations
@@
$\frac{1}{A}\sum_{\mathbf{k},\mathbf{k}^{\prime}}V_{DW}(-\mathbf{k}-\mathbf{k}^{\prime}+\mathbf{Q})f^{\dagger}_{c,\mathbf{k}}f^{\dagger}_{v,-\mathbf{k}+\mathbf{Q}}f_{v,-\mathbf{k}^{\prime}+\mathbf{Q}}f_{c,\mathbf{k}^{\prime}}+V_{DW}(-\mathbf{k}-\mathbf{k}^{\prime}-\mathbf{Q})f^{\dagger}_{c,\mathbf{k}}f^{\dagger}_{v,-\mathbf{k}-\mathbf{Q}}f_{v,-\mathbf{k}^{\prime}-\mathbf{Q}}f_{c,\mathbf{k}^{\prime}}$ (40)
@@
$\Delta_{\pm,n}(\mathbf{k})$ $=\frac{1}{A}\sum_{\mathbf{k}^{\prime},m}V_{DW}(-\mathbf{k}-\mathbf{k}^{\prime}\pm(1-2n)\mathbf{Q}-2m\mathbf{Q})\langle f_{v,-\mathbf{k}^{\prime}\pm\mathbf{Q}-2m\mathbf{Q}}f_{c,\mathbf{k}^{\prime}+2m\mathbf{Q}}\rangle.$ (41)
@@
$H_{\text{BdG}}=\sum_{\mathbf{k}}\begin{pmatrix}f^{\dagger}_{c,\mathbf{k}+2\mathbf{Q}}&f_{v,-\mathbf{k}+\mathbf{Q}}&f^{\dagger}_{c,\mathbf{k}}&f_{v,-\mathbf{k}-\mathbf{Q}}&f^{\dagger}_{c,\mathbf{k}-2\mathbf{Q}}\end{pmatrix}\begin{pmatrix}\xi_{c,\mathbf{k}-2\mathbf{Q}}&0&0&\Delta_{+,2}(\mathbf{k})&0\\ 0&-\xi_{v,-\mathbf{k}+\mathbf{Q}}&\Delta_{+,1}(\mathbf{k})&0&\Delta_{-,2}(\mathbf{k})\\ 0&\Delta^{*}_{+,1}(\mathbf{k})&\xi_{c,\mathbf{k}}&\Delta_{-,1}(\mathbf{k})&0\\ \Delta^{*}_{+,2}(\mathbf{k})&0&\Delta^{*}_{-,1}(\mathbf{k})&-\xi_{v,-\mathbf{k}-\mathbf{Q}}&0\\ 0&\Delta^{*}_{-,2}(\mathbf{k})&0&0&\xi_{c,\mathbf{k}+2\mathbf{Q}}\end{pmatrix}\begin{pmatrix}f_{c,\mathbf{k}+2\mathbf{Q}}\\ f^{\dagger}_{v,-\mathbf{k}+\mathbf{Q}}\\ f_{c,\mathbf{k}}\\ f^{\dagger}_{v,-\mathbf{k}-\mathbf{Q}}\\ f_{c,\mathbf{k}-2\mathbf{Q}}\end{pmatrix}.$ (42)
We now numerically diagonalize $H_{\text{BdG}}$ and solve for the gap functions self-consistently. Fig. 11 plots the obtained $\Delta_{\pm,1}$ as a function of $\mathbf{k}$.
Figure 11: Quantum critical fluctuation-mediated superconductivity. Pair-density wave (PDW) gap for $n=-0.36\times 10^{12}$ cm-2 and $V_{b}=4.2$ meV obtained using a $79\times 79$ momentum mesh.
@@
$\frac{1}{A}\sum_{\mathbf{k},\mathbf{k}^{\prime}}V_{DW}(-\mathbf{k}-\mathbf{k}^{\prime})f^{\dagger}_{c,\mathbf{k}}f^{\dagger}_{v,-\mathbf{k}}f_{v,-\mathbf{k}^{\prime}}f_{c,\mathbf{k}^{\prime}}.$ (43)
@@ Appendix F NFL Properties
$\Pi_{cv}(\mathbf{q},\Omega)=\int\frac{d^{2}kd\omega}{(2\pi)^{3}}\frac{1}{i(\omega+\Omega)-v_{c}(k_{x}+q_{x})-\frac{(k_{y}+q_{y})^{2}}{2m_{c}}}\frac{1}{i\omega+v_{v}k_{x}+\frac{k_{y}^{2}}{2m_{v}}}.$ (44)
Low-energy electron scatterings correspond to $q_{x}=0$ and $q_{y}\neq 0$, therefore, we set $q_{x}=0$, in order to extract the leading singular contribution. Let us also assume $m_{c}=m_{v}=m$ , which will not affect our conclusions qualitatively [^2] We generically consider $v_{c}\neq v_{v}$. For the case $v_{c}=v_{v}$, we note that $\Pi_{cv}$ exhibits an logarithmic IR divergence in the static limit within our model: $\Omega(\mathbf{q},\Omega=0)\sim\log|q_{y}|$..
@@
$\Pi_{cv}(\mathbf{q},\Omega)=-i\int\frac{dk_{y}d\omega}{(2\pi)^{2}}\frac{\theta(\omega+\Omega)-\theta(-\omega)}{i((v_{v}+v_{c})\omega+v_{v}\Omega)+v_{c}\frac{k_{y}^{2}}{2m}-v_{v}\frac{(k_{y}+q_{y})^{2}}{2m}}.$ (45)
@@
$\Pi_{cv}(\mathbf{q},\Omega)=\frac{1}{2\pi(v_{c}+v_{v})}\int\frac{dk_{y}}{2\pi}\log\left(-i\Omega v_{c}+v_{c}\frac{k_{y}^{2}}{2m}-v_{v}\frac{(k_{y}+q_{y})^{2}}{2m}\right)+\log\left(i\Omega v_{v}+v_{c}\frac{k_{y}^{2}}{2m}-v_{v}\frac{(k_{y}+q_{y})^{2}}{2m}\right)$ (46)
@@
$\Pi_{cv}(\mathbf{q},\Omega)\approx\frac{m}{2\pi(v_{c}+v_{v})}\left(\sqrt{\frac{v_{c}}{v_{v}}}+\sqrt{\frac{v_{v}}{v_{c}}}\right)\frac{|\Omega|}{|q_{y}|},$ (47)
@@
$S_{b}=\sum_{s}\int\frac{d^{2}qd\Omega}{(2\pi)^{3}}\left(\Omega^{2}+(\mathbf{q}+s\mathbf{Q})^{2}+\frac{mg^{2}}{2\pi(v_{c}+v_{v})}\left(\sqrt{\frac{v_{c}}{v_{v}}}+\sqrt{\frac{v_{v}}{v_{c}}}\right)\frac{|\Omega|}{|q_{y}|}\right)|b_{s}(\mathbf{q},\Omega)|^{2}.$ (48)
This yields a dynamical critical exponent of $z=3$, which leads to specific heat scaling at low temperatures as $C\sim T^{2/3}$ Lee (2018). Using the modified boson propagator, we calculate the one-loop fermion self-energy. The valence electron self-energy is,
$\Sigma_{v}(\mathbf{k},\omega)=g^{2}\int\frac{d^{2}qd\Omega}{(2\pi)^{3}}\frac{1}{q_{y}^{2}+\frac{mg^{2}}{2\pi(v_{c}+v_{v})}\left(\sqrt{\frac{v_{c}}{v_{v}}}+\sqrt{\frac{v_{v}}{v_{c}}}\right)\frac{|\Omega|}{|q_{y}|}}G_{c}(\mathbf{k}+\mathbf{q},\omega+\Omega),$ (49)
@@
$\Sigma_{v}(\omega)\sim i|\omega|^{2/3}\text{sgn}(\omega)$ (50)
indicating a breakdown of Fermi liquid quasiparticles near the hotspots. We comment that the self-energies obtained here are qualitatively the same as the problem of a Fermi surface coupled to a $\mathbf{Q}=0$ critical boson order parameter Sachdev (2011). In this case, the Landau quasiparticles get destabilized everywhere on the Fermi surface, but in contrast, in our case, only the fermions near the hot-spots are affected. Note also that the answers obtained here are distinct from the commonly analyzed case of spin-density wave criticality in cuprates, where the boson self-energy is linear in $|\Omega|$ with a constant $q$-independent denominator. This further leads to deviations in the fermionic self-energy as well. The main qualitative difference is that the Fermi velocities at the hot-spots connected by the ordering wave-vector are non-collinear in the SDW case, whereas here, they are collinear.
Motivated by previous work Senthil (2014) on diverging quantum oscillation mass near a quantum phase transition, we analyze this possibility in the present context. The quantum oscillation mass $m_{\text{QO}}$ is expressed as a line integral over the Fermi surface: $m_{\text{QO}}=\frac{1}{2\pi}\oint\frac{dK}{v_{F}(\theta)}$, where $\theta$ labels Fermi surface patches, and $v_{F}(\theta)$ is the Fermi velocity. Let us approach the phase transition to the X-DW phase from the electron-hole plasma state where there is no exciton ordering. Here, we expect a Fermi liquid form for the fermion Green’s function:
$G(\mathbf{k},\omega)=\frac{Z}{i\omega-v_{F}(\theta)k_{\perp}}.$ (51)
[…]

### v2

@@ IV.2 Spiral X-DW
$H_{\delta\chi-f}$ $=-\sum_{\mathbf{q},\mathbf{k}}\delta\chi_{-\mathbf{q}}f_{c,\mathbf{k}}^{{\dagger}}f_{v,\mathbf{k}+\textbf{Q}+\mathbf{q}}+\text{h.c.}$
@@
$=-\sum_{\mathbf{q},\mathbf{k}}\vec{f}_{\mathbf{q},\mathbf{k}}^{\ {\dagger}}\Big[\vec{\delta\chi}({-\mathbf{q}})\cdot\vec{\sigma}\Big]\vec{f}_{\mathbf{q},\mathbf{k}},$ (A.28)
where we have used the fact that $\delta\chi_{r}=\delta\chi_{x}(\mathbf{r})+i\delta\chi_{y}(\mathbf{r})$; this implies that $\delta\chi_{x/y}(\mathbf{q})]^{*}=\delta\chi_{x/y}(-\mathbf{q})$, and $S_{x}(\mathbf{q})\equiv\sum_{\mathbf{k}}\left(f^{\dagger}_{v,\mathbf{k}+\mathbf{q}+\mathbf{Q}}f_{c,\mathbf{k}}+f^{\dagger}_{c,\mathbf{k}}f_{v,\mathbf{k}+\mathbf{q}+\mathbf{Q}}\right)$, $S_{y}(\mathbf{q})\equiv-i\sum_{\mathbf{k}}\left(-f^{\dagger}_{v,\mathbf{k}+\mathbf{q}+\mathbf{Q}}f_{c,\mathbf{k}}+f^{\dagger}_{c,\mathbf{k}}f_{v,\mathbf{k}+\mathbf{q}+\mathbf{Q}}\right)$, and $\vec{f}_{\mathbf{q},\mathbf{k}}=(f_{c,\mathbf{k}}\ f_{v,\mathbf{k}+\mathbf{q}+\mathbf{Q}})^{\text{T}}$ where we express the fermionic operators in the basis of the mean-field Hamiltonian Eq. A.10.
@@
$H_{fluc.}=\frac{A}{2}\sum_{\mathbf{q}}\delta\chi_{i}(-\mathbf{q})\left(\frac{2}{U_{0}}\delta_{ij}-\Pi_{ij}(\mathbf{q},\Omega=0)\right)\delta\chi_{j}(\mathbf{q}).$ (A.29)
@@
$\Pi_{ij}(\mathbf{q},\Omega)=-\int\frac{d^{2}kd\omega}{(2\pi)^{3}}\text{Tr}\left(\sigma_{i}G(\mathbf{k}+\mathbf{q},\omega+\Omega)\sigma_{j}G(\mathbf{k},\omega)\right)$ (A.30)
@@
$H_{fluc.}=\sum_{\mathbf{q}}\begin{pmatrix}\delta\chi_{x}(-\mathbf{q})&\delta\chi_{y}(-\mathbf{q})\end{pmatrix}\begin{pmatrix}\rho^{x}_{xx}q_{x}^{2}+\rho^{y}_{xx}q_{y}^{2}+m_{xx}&is_{xy}q_{x}\\ -is_{xy}q_{x}&\rho^{x}_{yy}q_{x}^{2}+\rho^{y}_{yy}q_{y}^{2}\end{pmatrix}\begin{pmatrix}\delta\chi_{x}(\mathbf{q})\\ \delta\chi_{y}(\mathbf{q})\end{pmatrix}$ (A.33)
$\rho_{xx/yy}$, $s_{xy}$ and $m_{xx}$ are fit parameters obtained numerically. The form of Eq. A.33 follows from the even (odd)-in-momentum property of the diagonal (off-diagonal) elements of $\Pi_{\alpha\beta}(\mathbf{q})$.
@@
$\sin X(\mathbf{q})$ $=\frac{-s_{xy}q_{x}}{\sqrt{\left((\rho^{x}_{xx}-\rho^{x}_{yy})q_{x}^{2}+(\rho^{y}_{xx}-\rho^{y}_{yy})q_{y}^{2}+m_{xx}\right)^{2}/4+s_{xy}^{2}q_{x}^{2}}}$ (A.34)
Our next step is to calculate the interactions between the collective modes and the low-energy fermionic quasiparticles, as done for the collear case. We shall focus on the Goldstone fluctuations in the following, since the gapped collective mode would mediate relatively weaker interactions between the fermions away from the quantum critical line. Since we are considering hole doping, the Fermi surface is formed by the lower band $f_{-}$, which is obtained by diagonalizing the mean-field Hamiltonian A.10 within the Hubbard approximation to yield eigenmodes,
$f_{-,\mathbf{k}}=-\sin\left(\frac{\theta(\mathbf{k})}{2}\right)f_{c,\mathbf{k}}+\cos\left(\frac{\theta(\mathbf{k})}{2}\right)f_{v,\mathbf{k}+\mathbf{Q}},$ (A.35)
where $\sin\theta(\mathbf{k})=-\frac{\chi}{E_{\mathbf{k}}}$. We can thus rewrite Eq. A.28 in terms of the Goldstone mode and lower-band fermion operators to obtain the spin-independent) fermion-Goldstone mode interaction $\alpha$, where the form factor is:
$\alpha(\mathbf{k}+\mathbf{q},\mathbf{k})=\cos\left(\frac{X(\mathbf{q})}{2}\right)\sin\left(\frac{\theta(\mathbf{k}+\mathbf{q})-\theta(\mathbf{k})}{2}\right)+\sin\left(\frac{X(\mathbf{q})}{2}\right)\sin\left(\frac{\theta(\mathbf{k}+\mathbf{q})+\theta(\mathbf{k})}{2}\right).$ (A.36)
Finally, integrating out the Goldstone mode, we obtain an effective four-fermion interaction term identical to the one derived for the collinear case in Eq.A.22. Fig.A.5 (right panel) plots the interaction $V_{G}$ as $l_{x}\rightarrow k_{x}$, $l_{y}=k_{y}$. The interaction is attractive in a region of momentum space. However, since the spiral X-DW state breaks $C^{\prime}_{2z}$, states at momenta $\mathbf{k}$ and $\mathbf{k}$ are not degenerate, and it would require strong attractive interactions for them to overcome the energy difference and pair. While such a superconducting instability can occur in principle, our mean-field calculations for superconductivity do not yield superconducting solutions in the parameter space explored in this work.
Figure A.5: Spiral X-DW. Left panel: Fermi surface of the spiral X-DW with yellow marking occupied states. Right panel: Goldstone mode mediated interaction matrix elements. Dashed lines plot the locus of points satisfying $\xi_{c,\mathbf{k}}=\xi_{v,\mathbf{k}\pm\mathbf{Q}}$ (curve to the left/right). The parameters used are $n=-0.7\times 10^{12}$ cm-2, $V_{b}=14.4$ meV and a $59\times 59$ momentum mesh.
V Superconductivity Mediated by Critical Fluctuations
@@
$\frac{1}{A}\sum_{\mathbf{k},\mathbf{k}^{\prime}}V_{DW}(-\mathbf{k}-\mathbf{k}^{\prime}+\mathbf{Q})f^{\dagger}_{c,\mathbf{k}}f^{\dagger}_{v,-\mathbf{k}+\mathbf{Q}}f_{v,-\mathbf{k}^{\prime}+\mathbf{Q}}f_{c,\mathbf{k}^{\prime}}+V_{DW}(-\mathbf{k}-\mathbf{k}^{\prime}-\mathbf{Q})f^{\dagger}_{c,\mathbf{k}}f^{\dagger}_{v,-\mathbf{k}-\mathbf{Q}}f_{v,-\mathbf{k}^{\prime}-\mathbf{Q}}f_{c,\mathbf{k}^{\prime}}$ (A.37)
@@
$\Delta_{\pm,n}(\mathbf{k})$ $=\frac{1}{A}\sum_{\mathbf{k}^{\prime},m}V_{DW}(-\mathbf{k}-\mathbf{k}^{\prime}\pm(1-2n)\mathbf{Q}-2m\mathbf{Q})\langle f_{v,-\mathbf{k}^{\prime}\pm\mathbf{Q}-2m\mathbf{Q}}f_{c,\mathbf{k}^{\prime}+2m\mathbf{Q}}\rangle.$ (A.38)
@@
$H_{\text{BdG}}=\sum_{\mathbf{k}}\begin{pmatrix}f^{\dagger}_{c,\mathbf{k}+2\mathbf{Q}}&f_{v,-\mathbf{k}+\mathbf{Q}}&f^{\dagger}_{c,\mathbf{k}}&f_{v,-\mathbf{k}-\mathbf{Q}}&f^{\dagger}_{c,\mathbf{k}-2\mathbf{Q}}\end{pmatrix}\begin{pmatrix}\xi_{c,\mathbf{k}-2\mathbf{Q}}&0&0&\Delta_{+,2}(\mathbf{k})&0\\ 0&-\xi_{v,-\mathbf{k}+\mathbf{Q}}&\Delta_{+,1}(\mathbf{k})&0&\Delta_{-,2}(\mathbf{k})\\ 0&\Delta^{*}_{+,1}(\mathbf{k})&\xi_{c,\mathbf{k}}&\Delta_{-,1}(\mathbf{k})&0\\ \Delta^{*}_{+,2}(\mathbf{k})&0&\Delta^{*}_{-,1}(\mathbf{k})&-\xi_{v,-\mathbf{k}-\mathbf{Q}}&0\\ 0&\Delta^{*}_{-,2}(\mathbf{k})&0&0&\xi_{c,\mathbf{k}+2\mathbf{Q}}\end{pmatrix}\begin{pmatrix}f_{c,\mathbf{k}+2\mathbf{Q}}\\ f^{\dagger}_{v,-\mathbf{k}+\mathbf{Q}}\\ f_{c,\mathbf{k}}\\ f^{\dagger}_{v,-\mathbf{k}-\mathbf{Q}}\\ f_{c,\mathbf{k}-2\mathbf{Q}}\end{pmatrix}$. (A.39)
We now numerically diagonalize $H_{\text{BdG}}$ and solve for the gap functions self-consistently. For our calculations, we truncate to keep the five lowest energy bands and obtain $\Delta_{+}=\Delta_{-}$. Fig. A.6 plots the obtained $\Delta_{\pm,1}$ as a function of $\mathbf{k}$.
Figure A.6: Quantum critical fluctuation-mediated superconductivity. Pair-density wave (PDW) gap for $n=-0.36\times 10^{12}$ cm-2 and $V_{b}=4.2$ meV obtained using a $79\times 79$ momentum mesh.
@@
$\frac{1}{A}\sum_{\mathbf{k},\mathbf{k}^{\prime}}V_{DW}(-\mathbf{k}-\mathbf{k}^{\prime})f^{\dagger}_{c,\mathbf{k}}f^{\dagger}_{v,-\mathbf{k}}f_{v,-\mathbf{k}^{\prime}}f_{c,\mathbf{k}^{\prime}}.$ (A.40)
@@ VI NFL Properties
$\Pi_{cv}(\mathbf{q},\Omega)=\int\frac{d^{2}kd\omega}{(2\pi)^{3}}\frac{1}{i(\omega+\Omega)-v_{c}(k_{x}+q_{x})-\frac{(k_{y}+q_{y})^{2}}{2m_{c}}}\frac{1}{i\omega+v_{v}k_{x}+\frac{k_{y}^{2}}{2m_{v}}}.$ (A.41)
Low-energy electron scatterings correspond to $q_{x}=0$ and $q_{y}\neq 0$, therefore, we set $q_{x}=0$, in order to extract the leading singular contribution. Let us also assume $m_{c}=m_{v}=m$ , which will not affect our conclusions qualitatively [^1] We generically consider $v_{c}\neq v_{v}$. For the case $v_{c}=v_{v}$, we note that $\Pi_{cv}$ exhibits a logarithmic IR divergence in the static limit within our model: $\Omega(\mathbf{q},\Omega=0)\sim\log|q_{y}|$..
@@
$\Pi_{cv}(\mathbf{q},\Omega)=-i\int\frac{dk_{y}d\omega}{(2\pi)^{2}}\frac{\theta(\omega+\Omega)-\theta(-\omega)}{i((v_{v}+v_{c})\omega+v_{v}\Omega)+v_{c}\frac{k_{y}^{2}}{2m}-v_{v}\frac{(k_{y}+q_{y})^{2}}{2m}}.$ (A.42)
@@
$\Pi_{cv}(\mathbf{q},\Omega)=\frac{1}{2\pi(v_{c}+v_{v})}\int\frac{dk_{y}}{2\pi}\log\left(-i\Omega v_{c}+v_{c}\frac{k_{y}^{2}}{2m}-v_{v}\frac{(k_{y}+q_{y})^{2}}{2m}\right)+\log\left(i\Omega v_{v}+v_{c}\frac{k_{y}^{2}}{2m}-v_{v}\frac{(k_{y}+q_{y})^{2}}{2m}\right)$ (A.43)
@@
$\Pi_{cv}(\mathbf{q},\Omega)\approx\frac{m}{2\pi(v_{c}+v_{v})}\left(\sqrt{\frac{v_{c}}{v_{v}}}+\sqrt{\frac{v_{v}}{v_{c}}}\right)\frac{|\Omega|}{|q_{y}|},$ (A.44)
@@
$S_{b}=\sum_{s}\int\frac{d^{2}qd\Omega}{(2\pi)^{3}}\left(\Omega^{2}+(\mathbf{q}+s\mathbf{Q})^{2}+\frac{mg^{2}}{2\pi(v_{c}+v_{v})}\left(\sqrt{\frac{v_{c}}{v_{v}}}+\sqrt{\frac{v_{v}}{v_{c}}}\right)\frac{|\Omega|}{|q_{y}|}\right)|b_{s}(\mathbf{q},\Omega)|^{2}.$ (A.45)
This yields a dynamical critical exponent of $z=3$, which leads to specific heat scaling at low temperatures as $C\sim T^{2/3}$ [31]. Using the modified boson propagator, we calculate the one-loop fermion self-energy. The valence electron self-energy is,
$\Sigma_{v}(\mathbf{k},\omega)=g^{2}\int\frac{d^{2}qd\Omega}{(2\pi)^{3}}\frac{1}{q_{y}^{2}+\frac{mg^{2}}{2\pi(v_{c}+v_{v})}\left(\sqrt{\frac{v_{c}}{v_{v}}}+\sqrt{\frac{v_{v}}{v_{c}}}\right)\frac{|\Omega|}{|q_{y}|}}G_{c}(\mathbf{k}+\mathbf{q},\omega+\Omega),$ (A.46)
@@
$\Sigma_{v}(\omega)\sim i|\omega|^{2/3}\text{sgn}(\omega)$ (A.47)
indicating a breakdown of Fermi liquid quasiparticles near the hotspots. We comment that the self-energies obtained here are qualitatively the same as the problem of a Fermi surface coupled to a $\mathbf{Q}=0$ critical boson order parameter [54]. In this case, the Landau quasiparticles get destabilized everywhere on the Fermi surface, but in contrast, in our case, only the fermions near the hot-spots are affected. Note also that the answers obtained here are distinct from the commonly analyzed case of spin-density wave criticality in cuprates, where the boson self-energy is linear in $|\Omega|$ with a constant $q$-independent denominator. This further leads to deviations in the fermionic self-energy as well. The main qualitative difference is that the Fermi velocities at the hot-spots connected by the ordering wave-vector are non-collinear in the SDW case, whereas here, they are collinear.
Motivated by previous work [56] on diverging quantum oscillation mass near a quantum phase transition, we analyze this possibility in the present context. The quantum oscillation mass $m_{\text{QO}}$ is expressed as a line integral over the Fermi surface: $m_{\text{QO}}=\frac{1}{2\pi}\oint\frac{dK}{v_{F}(\theta)}$, where $\theta$ labels Fermi surface patches, and $v_{F}(\theta)$ is the Fermi velocity. Let us approach the phase transition to the X-DW phase from the electron-hole plasma state where there is no exciton ordering. Here, we expect a Fermi liquid form for the fermion Green’s function:
$G(\mathbf{k},\omega)=\frac{Z}{i\omega-v_{F}(\theta)k_{\perp}}.$ (A.48)
Near the hot-spots and close to criticality, we anticipate the following scaling form for the Green’s function: $G\sim g(\frac{\omega}{k_{\parallel}^{z}},\frac{k_{\perp}}{k_{\parallel}^{2}},\frac{\omega}{\delta})$, where $\delta$ is an energy scale that characterizes the distance to the quantum critical point, $k_{\perp}$ ($k_{\parallel}$) quantifies the momentum deviation from the hotspots along the normal (tangential) direction to the Fermi surface. Comparing with the Fermi liquid form in Eq. A.48, we obtain $v_{F}\sim|k_{\parallel}|^{z-2}$. Since $z=3$ in our case (on approach to the quantum critical point) we expect that $v_{F}\sim|k_{\parallel}|$ and thus vanishes at the hot spot. Therefore, $m_{\text{QO}}\sim\oint dk_{\parallel}/k_{\parallel}$ diverges logarithmically.